OpenStudy (anonymous):

A box of rotini noodles contains 10 servings. If you eat 7/8 of the recommended serving as a meal, how many complete meals can you get from one box?

OpenStudy (anonymous):

You've eaten 7/8, which is how much over 10? If I'm interpreting the question correctly, that's what you need to work through. Then since they're looking for a complete meal, you'll have to round the number out of ten to the nearest whole number--always rounding down. If we let x be the amount remaining of 10 servings, \[\frac{ 7 }{ 8 }=\frac{ x }{ 10 }\]

OpenStudy (anonymous):

Solve for x, round down, and you have your answer.

OpenStudy (anonymous):

i can't figure it out

OpenStudy (anonymous):

You cross-multiple, or multiple both sides by 8 and by 10. \[\frac{ 7 }{ 8 }=\frac{ x }{ 10 }\]\[(7)\times(10)=(x)\times(8)\]\[70=8x\]\[\frac{ 70 }{ 8 }=\frac{ 8x }{ 8 }\]\[\frac{ 70 }{8 }=x\]Punch that into a calculator and you'll get an answer, just remember to round down because we're looking for full meals.